OpenStudy (anonymous):

|3 - x| = 13 Can somebody explain how to do this? I think i know how but im not sure.

OpenStudy (anonymous):

3-x=13 and 3-x=-13 -x=13-3 and -x=-13-3 -x=10 and -x=-16 x=-10 and x=16

OpenStudy (helder_edwin):

remeber the theorem \[ |a|=b \Leftrightarrow a=\pm b \]

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

the complete theorem says \[ |a|=b \Leftrightarrow \begin{cases} b\geq0\\ \wedge\\ a=\pm b \end{cases} \] in your case b=13>0

OpenStudy (anonymous):

thanks